Other Applications

• exp, sin, cos, log, arctan, etc.

• π (Chudnovsky’s formula, a Ramanujan type series)

• hypergeometric functions at x ∈ ℚ

• hypergeometric and holonomic functions at x ∈ ℂ [J. van der Hoeven]

• Γ(x) at x ∈ ℚ

• Large terms of P-recursive sequences, i.e. f_{n} (n large) where f(x) = ∑ f_{n}x^{n} is holonomic,

• ζ(k), k > 1 odd (using Cohen-Villegas-Zagier convergence acceleration)